# Self-Supervised Multi-Part Articulated Objects Modeling via Deformable Gaussian Splatting and Progressive Primitive Segmentation

### v1

@@ 5.1 Experimental Setting
Datasets.
Most studies evaluate only one object per category, which limits the statistical power. We extend PartNet-Mobility [1] by selecting eight categories and two to three instances per category, each with two movable parts, so all methods face the same kinematic complexity. We also include four objects with three movable parts to test the robustness under higher joint counts. Real-scene performance is measured on the RS-Art we proposed. Together, the enlarged PartNet-Mobility split and RS-Art form a balanced test bed for synthetic and real scenarios.
Baselines.
We compare our method with recent SOTA methods. The single-joint baselines, which can process only one movable part at a time, are PARIS [20] and ArticulatedGS [21]. As ArticulatedGS [21] has no public code, we reproduced it with the official 3DGS implementation and the details described in the article. Each single-joint pipeline was run once for every joint: in each run, the input images show motion of a single joint while the other parts remain fixed, and the per-joint outputs were merged for evaluation. The multi-joint baselines, which can handle several movable parts in one pass, are DTArt [23] and ArtGS [24]. Both baselines require a prespecified part count, whereas our method infers the parts automatically.
Metrics.
We adopt exactly the evaluation metrics defined in DTArt [23], covering the geometry, part segmentation and precision of the joint parameter. The only extension is that our reported scores are aggregated over all movable parts of an object, rather than being limited to a single target part.
@@ 5.2 Comparison on Objects with Multiple Movable Parts
Tab. 1 reports quantitative results on our expanded PartNet-Mobility split, which contains objects with two independently actuated parts. Previous studies usually average across several runs, yet we observe large run-to-run variance in several baselines, to reveal this sensitivity, we list the first run for every method. These numbers are not cherry-picked, and subsequent runs often produce even lower scores. DeGSS needs no manual part specification and still outperforms the baselines in most metrics. Fig. 2, which also shows an object with three movable parts, qualitatively confirms that our joint estimates align more closely with the ground truth. DTArt [23] benefits from depth supervision and therefore renders smoother surfaces, but still lags behind in articulation accuracy. Overall, the use of a latent-conditioned deformation field spanning multiple interaction states yields markedly higher stability than single-state or pairwise pipelines. Results on more objects with three movable parts are presented in Sec. D.1 of the Appendix, where the same trend persists.
@@ 5.3 Generalization to Unseen Interaction States
Fig. 3 illustrates that our method is the only one that is capable of disentangling and controlling the motion of several parts independently while maintaining plausible geometry in interaction states never shown during training. By interpolating the latent code and deforming only the Gaussians that belong to a chosen part, we generate smooth, collision-free trajectories for every drawer and for the door. In contrast, PARIS [20] cannot keep the parts separated and often collapses or blends them together. The capacity to vary multiple joints independently and to synthesize genuinely novel poses represents a substantial step toward high-fidelity digital twin modeling of articulated objects.
@@ 5.4 Result on the RS-Art Dataset
Figure 4: Real object validation on RS-Art.
@@ 5.5 Ablation Studies
The most informative ablation is to remove the coarse-to-fine refinement stage introduced in Sec. 3.3 and Sec. 3.4. Fig. 5 contrasts the results without of the refinement stage (top) with those with of it (bottom) in a cabinet and fridge that contain multiple movable panels. Without refinement, the reconstructed parts bleed into one another and the contact regions exhibit severe artifacts; after refinement, the mesh shows clean separation, sharper boundaries, and no interpenetration. The improvement stems from visibility-guided SAM prompting and boundary-aware Gaussian splitting, which force the Gaussian field to conform to the per-part masks. We observe the same trend across other objects and provide additional visual and quantitative comparisons in Sec. D.3 of the Appendix.
Figure 5: Ablation on the coarse–to–fine refinement stage. For the variant without refinement we reconstruct multiple parts directly from the K-means clusters obtained in the coarse stage.
@@ 6 Conclusion
We introduce DeGSS, the fully self-supervised framework for articulated object modeling within a unified 3D Gaussian Splatting paradigm. A latent-conditioned deformation network warps a standard Gaussian field across interaction states, and a visibility-guided coarse-to-fine scheme refines deformation trajectories into part-aware Gaussian clusters. The resulting field supports part-level mesh extraction, joint typing, and motion parameter estimation. In an expanded PartNet-Mobility split and the novel RS-Art dataset, DeGSS surpasses existing methods without requiring part count in advance, and its learned deformation space generalizes smoothly to unseen interaction states, a property well suited to digital twin scenarios. The present implementation assumes accurate camera poses and cannot reconstruct regions that remain fully occluded in all views, leaving robustness to pose noise and reasoning about unseen structures as directions for future work.
@@ A.1 Boundary-Aware Gaussian Splitting
Selecting boundary candidates.
For each 3D Gaussian primitive $g_{i}=(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$, we first locate its ray–sampling nearest view $v_{\text{near}}$. Assuming a local affine projection, the covariance $\boldsymbol{\Sigma}_{i}$ is mapped to the image plane by
$\boldsymbol{\Sigma}^{\prime}_{i}\;=\;\mathbf{J}\mathbf{W}\boldsymbol{\Sigma}_{i}\mathbf{W}^{\!\top}\mathbf{J}^{\!\top},$ (13)
where $\mathbf{W}$ and $\mathbf{J}$ denote the camera projection matrix and its Jacobian, respectively [19]. The resulting $2\times 2$ matrix $\boldsymbol{\Sigma}^{\prime}_{i}$ defines an ellipse; its major axis is aligned with the eigenvector of $\boldsymbol{\Sigma}^{\prime}_{i}$ that has the largest eigenvalue. If either end of this axis lies outside the part mask $\mathcal{M}^{(v_{\text{near}})}_{p}$, the primitive is marked as boundary candidate.
Geometric split ratio.
Let $A^{\ast}$ and $B^{\ast}$ be the two end points of the projected major axis and let $O^{\ast}$ be their intersection with the mask boundary ($A^{\ast}$ lies inside, $B^{\ast}$ outside). Following the split strategy of Hu et al. [36], the in-mask proportion in image space is
@@
Because the projection is affine locally, the same ratio applies in 3D. With $\mathbf{e}$ denoting the unit eigenvector of $\boldsymbol{\Sigma}_{i}$ that corresponds to its largest eigenvalue and $s_{i}$ the length of the associated axis, the two children are
  • Part child: $\boldsymbol{\mu}_{i}^{\text{part}}=\boldsymbol{\mu}_{i}+\tfrac{1-\lambda_{2\text{D}}}{2}\,s_{i}\mathbf{e},\;s_{i}^{\text{part}}=\lambda_{2\text{D}}\,s_{i};$
  • Background child: $\boldsymbol{\mu}_{i}^{\text{bg}}=\boldsymbol{\mu}_{i}-\tfrac{\lambda_{2\text{D}}}{2}\,s_{i}\mathbf{e},\;s_{i}^{\text{bg}}=(1-\lambda_{2\text{D}})\,s_{i}.$
@@
Recursive splitting and mask consistency.
The background child is re-evaluated in its own nearest view; if its projected ellipse still crosses another part mask, the split is applied recursively until every descendant ellipse is fully contained in a single mask. The recursion stops when the ellipse lies entirely inside its mask or the new scale falls below a minimum threshold. After each split, the children are queued in descending order of projected error, so the most inaccurate primitives are processed first. Finally, a multiview voting scheme assigns a consistent mask label to every new Gaussian, ensuring cross-view agreement.
@@ A.2 Mesh Extraction from the Part-Aware Gaussian Field
Given the part-aware Gaussian field $\mathcal{G}^{(\boldsymbol{\alpha})}=\{\,\mathcal{G}^{(\boldsymbol{\alpha})}_{p}\,\}_{p=1}^{n_{\text{parts}}}$ for the interaction state $\boldsymbol{\alpha}$, we recover explicit geometry by multiview depth rendering followed by implicit–to–explicit surface reconstruction. For each subset of parts $\mathcal{G}^{(\boldsymbol{\alpha})}_{p}$, we render a depth map in every calibrated view, treating each Gaussian primitive as an ellipsoid whose pose and covariance are projected onto the image plane using the standard 3DGS rasterizer. The rendered depths are fused with a truncated signed-distance function (TSDF); voxel size and the truncation threshold are chosen empirically to balance detail and noise. The resulting TSDF volume provides an implicit surface for part $p$, from which we extract a triangle mesh $\mathbf{M}_{p}\!\in\!\mathbb{R}^{|V_{p}|\times 3}$ using the Marching Cubes algorithm implemented in Open3D [42]. This procedure yields a high-resolution, watertight mesh for every rigid part, faithfully capturing its geometry while preserving the spatial correspondence with the underlying Gaussian field.
@@ A.3 Joint Type Identification
For a part $p$ we gather Gaussian centroids in two states $P=\{p_{i}\}$ and $Q=\{q_{i}\}$, and align them with the Kabsch algorithm [41], which yields rotation $\mathbf{R}$ and translation $\mathbf{t}$. The residuals $d_{i}=q_{i}-(\mathbf{R}p_{i}+\mathbf{t})$ are stacked in the $3\times n$ matrix $\mathbf{D}=[\,d_{1}\;\cdots\;d_{n}]$, where the $n$ is the number of Gaussian kernels.
@@ A.4 Joint Parameter Estimation
Revolute joint.
A revolute joint is parameterized by a pivot point $\mathbf{p}\!\in\!\mathbb{R}^{3}$ and a unit quaternion $\mathbf{q}\!\in\!\mathbb{R}^{4}$ with $\lVert\mathbf{q}\rVert=1$. From the Kabsch alignment [41] we obtain a rotation matrix $\mathbf{R}\!\in\!\mathbb{R}^{3\times 3}$. Its rotation axis is the normalized eigenvector $\mathbf{u}$ that satisfies $\mathbf{R}\mathbf{u}=\mathbf{u}$, and the rotation angle is $\theta=\arccos\bigl{(}(\operatorname{tr}(\mathbf{R})-1)/2\bigr{)}$. The pivot $\mathbf{p}$ is found by least-squares fitting the axis to the static centroids $\mathcal{C}_{\text{static}}$,
@@
Prismatic joint.
A prismatic joint is defined by a unit slide axis $\mathbf{a}\!\in\!\mathbb{R}^{3}$ and a translation distance $d$. Let $P=\{p_{i}\}$ and $Q=\{q_{i}\}$ be the part centroids in the two states. Their means are $\bar{p}=\tfrac{1}{n}\sum_{i}p_{i}$ and $\bar{q}=\tfrac{1}{n}\sum_{i}q_{i}$, so the rigid translation is $\mathbf{t}=\bar{q}-\bar{p}$. Hence $\mathbf{a}=\mathbf{t}/\lVert\mathbf{t}\rVert$ and $d=\lVert\mathbf{t}\rVert$. These parameters constitute the minimal motion model used in our evaluation.
@@ Appendix B RS-Art Dataset
Capture platform.
[…]
The operator mounts an Intel RealSense RGB-D camera [43] rigidly on the lid of a laptop, then walks a circular trajectory whose radius is chosen in real time from the live video feed to keep the target object at the desired image scale. Each circuit produces a set of synchronized RGB and depth frames that share a consistent chroma-key background, which simplifies subsequent background removal and calibration.
@@
Real-scene acquisition pipeline.
As illustrated in Fig. 7, to keep the real data fully compatible with the experimental settings used for the PartNet-Mobility dataset [1], we record two complementary interaction groups for each object. (i) Multi-part sequences: all movable parts are actuated simultaneously, and the capture spans four evenly spaced configurations that bracket the mechanical limits, for example, fully closed, half-open, three-quarter open, and fully open. (ii) Single-part sequences: each rotational or translational part is moved while the others remain fixed; for every part, we acquire two extreme poses corresponding to its minimum and maximum joint limits.
For each interaction state, we sample RGB-D images along a circular path whose viewing elevation is uniformly chosen in the range $30^{\circ}\!\sim\!60^{\circ}$. The operator walks once around the object with an Intel RealSense camera [43] rigidly mounted on a laptop lid, keeping the target centered in the live preview.
@@
Modeling and registration workflow.
All digital assets are created in Isaac Sim 4.2 [44], whose USD pipeline provides both photorealistic rendering and physically consistent simulation (see Fig. 8 for a representative modeling scene). We first reverse-model each real object with calipers, producing a watertight USD mesh; subparts are manually segmented according to functional affordances and joint boundaries. High-resolution textures and PBR materials are authored in Blender 4.2.0 [45] and imported into Isaac Sim without loss of visual fidelity. The physical properties - mass, inertia, and friction - are assigned following the official Omniverse workflow [46]. The joint pivots and motion limits are calibrated by system identification, ensuring that the articulation of the digital twin in Isaac Sim reproduces the joint behavior observed on the physical object. The resulting USD files provide a pixel-accurate appearance and dynamics-ready articulation, forming a reliable ground truth for both vision and physics experiments.
@@
Dataset structure.
Fig. 9 outlines the RS-Art hierarchy. The dataset contains 18 articulated objects drawn from six everyday categories:drawer, lamp, eyeglasses, floppy-disk drive, basket, and phone / laptop stand with three instances per class. For every object we supply the complete set of multi-view RGB-D captures for each recorded interaction state together with the shared camera intrinsics and the per-view extrinsics; a fully reverse-modeling digital twin comprising a USD file that bundles part meshes, textures, physics properties, and joint definitions, a matching lightweight URDF, and watertight PLY meshes for geometry evaluation; and a concise meta.json that enumerates all parts, joint types, axes, limits, and labels every state as either multi-part or single-part motion. This combination of real imagery and fully articulated digital twins enables training, quantitative evaluation, and sim-to-real studies for part-aware geometry reconstruction and joint parameter estimation.
Appendix C Experiments Settings

### v2

@@ 5.1 Experimental Setting
Datasets. Most studies evaluate only one object per category, which limits the statistical power. We extend PartNet-Mobility (Xiang et al., 2020) by selecting eight categories and two to three instances per category, each with two movable parts, so all methods face the same kinematic complexity. We also include four objects with three movable parts to test the robustness under higher joint counts. Real-scene performance is measured on the RS-Art we proposed. Together, the enlarged PartNet-Mobility split and RS-Art form a well-balanced platform for both synthetic and real-world scenarios.
Baselines. We compare our method with recent SOTA methods. The single-joint baselines include PARIS (Liu et al., 2023) and ArticulatedGS (Guo et al., 2025). Each single-joint pipeline was run once for every joint: in each run, the input images show motion of a single joint while the other parts remain fixed, and the per-joint outputs were merged for evaluation. The multi-joint baselines include DTArt (Weng et al., 2024) and ArtGS (Liu et al., 2025b). Both baselines require a prespecified part count, whereas our method infers the parts automatically. For the dataset setup, two-state baselines sample 100 views per state (200 RGB images; DTArt also requires depth). Matching this budget, we sample 50 views across four states, totaling 200 RGB inputs. For more information on our method restricted to two interaction states, refer to Sec. D.1 of the Appendix.
Metrics. We adopt exactly the evaluation metrics defined in DTArt (Weng et al., 2024), covering the geometry reconstruction and precision of the joint parameter. CD-w, CD-s, and CD-m are used to measure the Chamfer Distance for the whole object, static parts, and movable parts in mesh reconstruction. Axis Ang, Axis Pos, and Part Motion are used to evaluate the errors between predicted and ground-truth joint axes parameters. The only extension is that our reported scores are aggregated over all movable parts of an object, rather than being limited to a single target part.
@@ 5.2 Comparison on Objects with Multiple Movable Parts
Tab. 1 reports quantitative results on our expanded PartNet-Mobility split, which contains objects with two independently actuated parts. Previous studies usually average across several runs, yet we observe large run-to-run variance in several baselines, to reveal this sensitivity, we list the first run for every method. These numbers are not cherry-picked, and subsequent runs often produce even lower scores. PD2GS needs no manual part specification and still outperforms the baselines in most metrics. Fig. 3, which also shows an object with three movable parts, qualitatively confirms that our joint estimates align more closely with the ground truth. DTArt (Weng et al., 2024) benefits from depth input and therefore renders smoother surfaces, but still lags behind in articulation accuracy. Overall, the use of a latent-conditioned deformation field spanning multiple interaction states yields markedly higher stability than single-state or pairwise pipelines. Results on more objects with three movable parts are presented in Sec. D.2 of the Appendix, where the same trend persists.
@@
Figure 5: Qualitative reconstruction results on intermediate states.
@@ 5.3 Generalization to Unseen Interaction States
Fig. 4 and Fig. 5 illustrate that our method is the only one that is capable of disentangling and controlling the motion of several parts independently while maintaining plausible geometry in interaction states never shown during training. Fig. 5 specifically demonstrates how we perform interpolation between the start state and the end state. By interpolating the latent code and deforming only the Gaussians that belong to a chosen part, we generate smooth, collision-free trajectories for every drawer and for the door. In contrast, PARIS (Liu et al., 2023) cannot keep the parts separated, often leading to part overlap or blending due to the inherent multi-solution nature of its method. The capacity to vary multiple joints independently and to synthesize genuinely novel poses represents a substantial step toward high-fidelity digital twin modeling of articulated objects.
Table 2: Quantitative comparison with deformable NeRF- and 3DGS-based methods on unseen-state rendering.
[…]
D-NeRF (Pumarola et al., 2021) 18.38 0.86 0.19
D2NeRF (Wu et al., 2022) 21.34 0.85 0.16
Dynamic3DGS (Katsumata et al., 2023) 25.96 0.89 0.17
4D-GS (Wu et al., 2024) 23.61 0.92 0.13
ArtGS (Liu et al., 2025b) 26.82 0.92 0.12
PD2GS (Ours) 28.02 0.93 0.09
To further substantiate this capability, we extend our analysis by comparing our method with recent dynamic scene modeling approaches (Pumarola et al., 2021; Wu et al., 2022; Yang et al., 2024c; Luiten et al., 2024), including deformable NeRF- and 3DGS-based method. Following the protocol in Tab. 1, we evaluate on the Storage category, supplying each method with four discrete interaction states as input. We assess the capability of each method to synthesize intermediate interaction states by rendering novel views interpolated between input configurations. Additionally, for the two-state-based ArtGS (Liu et al., 2025b), we apply its learned joint parameters to rigidly transform the reconstructed Gaussian field to generate the corresponding interpolated poses. As shown in Fig. 6 and Tab. 2, our method achieves significantly superior performance in both visual fidelity and quantitative metrics. Notably, it effectively eliminates geometric artifacts that commonly plague holistic deformation methods.
Figure 6: Qualitative comparison with deformable NeRF- and 3DGS-based methods on unseen-state rendering.
Figure 7: Real object validation on RS-Art.
@@ 5.4 Result on the RS-Art Dataset
Figure 8: Ablation on the coarse–to–fine refinement stage.
@@ 5.5 Ablation Studies
To evaluate the effectiveness of our coarse-to-fine strategy (Sec. 3.3 and 3.4), we conduct an ablation study by removing the refinement stage. Fig. 8 compares the results without the refinement stage (top) to those with it (bottom) on an oven containing multiple movable panels. For the variant without refinement, we reconstruct multiple parts directly from the K-means clusters obtained in the coarse stage. Without refinement, the reconstructed parts bleed into one another and exhibit artifacts; after refinement, the mesh shows clean separation, sharper boundaries, and no interpenetration. The improvement stems from visibility-guided SAM prompting and boundary-aware Gaussian splitting, which force the Gaussian field to conform to the per-part masks. We provide additional visual and quantitative comparisons, along with further ablation analysis on movable-component count estimation via VLM, SAM-based segmentation, and each pipeline module, in Sec. D.7 of the Appendix.
@@ 6 Conclusion
We introduce PD2GS, the fully self-supervised framework for articulated object modeling within a unified 3D Gaussian Splatting paradigm. A latent-conditioned deformation network warps a canonical Gaussian field across interaction states, and a visibility-guided coarse-to-fine scheme refines deformation trajectories into part-aware Gaussian clusters. The resulting field supports part-level mesh extraction, joint typing, and motion parameter estimation. In an expanded PartNet-Mobility split and the novel RS-Art dataset, PD2GS surpasses existing methods without requiring part count in advance, and its learned deformation space generalizes smoothly to unseen interaction states, a property well suited to digital twin scenarios. The present implementation assumes accurate camera poses and cannot reconstruct regions that remain fully occluded in all views, leaving robustness to pose noise and reasoning about unseen structures as directions for future work.
Acknowledgements
This research was supported in part by the Natural Science Foundation of Anhui under Grant No. 2508085QF241, in part by the National Natural Science Foundation of China under Grant No. 62571055 and No. 62431015.
@@ A.1 Boundary-Aware Gaussian Splitting
Selecting boundary candidates. For each 3D Gaussian primitive $g_{i}=(\bm{\mu}_{i},\bm{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$, we first locate its ray–sampling nearest view $v_{\text{near}}$. Assuming a local affine projection, the covariance $\bm{\Sigma}_{i}$ is mapped to the image plane by
$\bm{\Sigma}^{\prime}_{i}\;=\;\mathbf{J}\mathbf{W}\bm{\Sigma}_{i}\mathbf{W}^{\!\top}\mathbf{J}^{\!\top},$ (13)
where $\mathbf{W}$ and $\mathbf{J}$ denote the camera projection matrix and its Jacobian, respectively (Kerbl et al., 2023). The resulting $2\times 2$ matrix $\bm{\Sigma}^{\prime}_{i}$ defines an ellipse; its major axis is aligned with the eigenvector of $\bm{\Sigma}^{\prime}_{i}$ that has the largest eigenvalue. If either end of this axis lies outside the part mask $\mathcal{M}^{(v_{\text{near}})}_{p}$, the primitive is marked as boundary candidate.
Geometric split ratio. Let $A^{\ast}$ and $B^{\ast}$ be the two end points of the projected major axis and let $O^{\ast}$ be their intersection with the mask boundary ($A^{\ast}$ lies inside, $B^{\ast}$ outside). Following the split strategy of Hu et al. (Hu et al., 2024), the in-mask proportion in image space is
@@
Because the projection is affine locally, the same ratio applies in 3D. With $\mathbf{e}$ denoting the unit eigenvector of $\bm{\Sigma}_{i}$ that corresponds to its largest eigenvalue and $s_{i}$ the length of the associated axis, the two children are
  • Part child: $\bm{\mu}_{i}^{\text{part}}=\bm{\mu}_{i}+\tfrac{1-\lambda_{2\text{D}}}{2}\,s_{i}\mathbf{e},\;s_{i}^{\text{part}}=\lambda_{2\text{D}}\,s_{i};$
  • Background child: $\bm{\mu}_{i}^{\text{bg}}=\bm{\mu}_{i}-\tfrac{\lambda_{2\text{D}}}{2}\,s_{i}\mathbf{e},\;s_{i}^{\text{bg}}=(1-\lambda_{2\text{D}})\,s_{i}.$
@@
Recursive splitting and mask consistency. The background child is re-evaluated in its own nearest view; if its projected ellipse still crosses another part mask, the split is applied recursively until every descendant ellipse is fully contained in a single mask. The recursion stops when the ellipse lies entirely inside its mask or the new scale falls below a minimum threshold. After each split, the children are queued in descending order of projected error, so the most inaccurate primitives are processed first. Finally, a multiview voting scheme assigns a consistent mask label to every new Gaussian, ensuring cross-view agreement.
@@ A.2 Mesh Extraction from the Part-Aware Gaussian Field
Given the part-aware Gaussian field $\mathcal{G}^{(\bm{\alpha})}=\{\,\mathcal{G}^{(\bm{\alpha})}_{p}\,\}_{p=1}^{n_{\text{parts}}}$ for the interaction state $\bm{\alpha}$, we recover explicit geometry by multiview depth rendering followed by implicit–to–explicit surface reconstruction. For each subset of parts $\mathcal{G}^{(\bm{\alpha})}_{p}$, we render a depth map in every calibrated view, treating each Gaussian primitive as an ellipsoid whose pose and covariance are projected onto the image plane using the standard 3DGS rasterizer. The rendered depths are fused with a truncated signed-distance function (TSDF); voxel size and the truncation threshold are chosen empirically to balance detail and noise. The resulting TSDF volume provides an implicit surface for part $p$, from which we extract a triangle mesh $\mathbf{M}_{p}\!\in\!\mathbb{R}^{|V_{p}|\times 3}$ using the Marching Cubes algorithm implemented in Open3D (Huang et al., 2024). This procedure yields a high-resolution, watertight mesh for every rigid part, faithfully capturing its geometry while preserving the spatial correspondence with the underlying Gaussian field.
@@ A.3 Joint Type Identification
For a part $p$ we gather Gaussian centroids in two states $P=\{p_{i}\}$ and $Q=\{q_{i}\}$, and align them with the Kabsch algorithm (Lawrence et al., 2019), which yields rotation $\mathbf{R}$ and translation $\mathbf{t}$. The residuals $d_{i}=q_{i}-(\mathbf{R}p_{i}+\mathbf{t})$ are stacked in the $3\times n$ matrix $\mathbf{D}=[\,d_{1}\;\cdots\;d_{n}]$, where the $n$ is the number of Gaussian kernels.
@@ A.4 Joint Parameter Estimation
Revolute joint. A revolute joint is parameterized by a pivot point $\mathbf{p}\!\in\!\mathbb{R}^{3}$ and a unit quaternion $\mathbf{q}\!\in\!\mathbb{R}^{4}$ with $\lVert\mathbf{q}\rVert=1$. From the Kabsch alignment (Lawrence et al., 2019) we obtain a rotation matrix $\mathbf{R}\!\in\!\mathbb{R}^{3\times 3}$. Its rotation axis is the normalized eigenvector $\mathbf{u}$ that satisfies $\mathbf{R}\mathbf{u}=\mathbf{u}$, and the rotation angle is $\theta=\arccos\bigl((\operatorname{tr}(\mathbf{R})-1)/2\bigr)$. The pivot $\mathbf{p}$ is found by least-squares fitting the axis to the static centroids $\mathcal{C}_{\text{static}}$,
@@
Prismatic joint. A prismatic joint is defined by a unit slide axis $\mathbf{a}\!\in\!\mathbb{R}^{3}$ and a translation distance $d$. Let $P=\{p_{i}\}$ and $Q=\{q_{i}\}$ be the part centroids in the two states. Their means are $\bar{p}=\tfrac{1}{n}\sum_{i}p_{i}$ and $\bar{q}=\tfrac{1}{n}\sum_{i}q_{i}$, so the rigid translation is $\mathbf{t}=\bar{q}-\bar{p}$. Hence $\mathbf{a}=\mathbf{t}/\lVert\mathbf{t}\rVert$ and $d=\lVert\mathbf{t}\rVert$. These parameters constitute the minimal motion model used in our evaluation.
@@ Appendix B RS-Art Dataset
Capture platform. Our data are acquired in a purpose-built chroma-key booth (Fig. 9). Four aluminum beams support a seamless green curtain that encloses a $2.5\,\text{m}$-high cubic volume and blocks stray reflections while maintaining uniform ambient illumination. In the center stands a square pedestal $0.5\,\text{m}\!\times\!0.5\,\text{m}$, fully wrapped in the same green fabric. The rear side of the cloth is glued to the floor so that the pedestal merges smoothly with the ground plane, eliminating visible seams in the depth and color images.
The operator mounts an Intel RealSense RGB-D camera (Keselman et al., 2017) rigidly on the lid of a laptop, then walks a circular trajectory whose radius is chosen in real time from the live video feed to keep the target object at the desired image scale. Each circuit produces a set of synchronized RGB and depth frames that share a consistent chroma-key background, which simplifies subsequent background removal and calibration.
@@
Real-scene acquisition pipeline. As illustrated in Fig. 10, to keep the real data fully compatible with the experimental settings used for the PartNet-Mobility dataset (Xiang et al., 2020), we record two complementary interaction groups for each object. (i) Multi-part sequences: all movable parts are actuated simultaneously, and the capture spans four evenly spaced configurations that bracket the mechanical limits, for example, fully closed, half-open, three-quarter open, and fully open. (ii) Single-part sequences: each rotational or translational part is moved while the others remain fixed; for every part, we acquire two extreme poses corresponding to its minimum and maximum joint limits.
For each interaction state, we sample RGB-D images along a circular path whose viewing elevation is uniformly chosen in the range $30^{\circ}\!\sim\!60^{\circ}$. The operator walks once around the object with an Intel RealSense camera (Keselman et al., 2017) rigidly mounted on a laptop lid, keeping the target centered in the live preview.
@@
Modeling and registration workflow. All digital assets are created in Isaac Sim 4.2 (NVIDIA, 2024b), whose USD pipeline provides both photorealistic rendering and physically consistent simulation (see Fig. 11 for a representative modeling scene). We first reverse-model each real object with calipers, producing a watertight USD mesh; subparts are manually segmented according to functional affordances and joint boundaries. High-resolution textures and PBR materials are authored in Blender 4.2.0 (Blender Foundation, 2024) and imported into Isaac Sim without loss of visual fidelity. The physical properties - mass, inertia, and friction - are assigned following the official Omniverse workflow (NVIDIA, 2024a). The joint pivots and motion limits are calibrated by system identification, ensuring that the articulation of the digital twin in Isaac Sim reproduces the joint behavior observed on the physical object. The resulting USD files provide a pixel-accurate appearance and dynamics-ready articulation, forming a reliable ground truth for both vision and physics experiments.
@@
Dataset structure. Fig. 12 outlines the RS-Art hierarchy. The dataset contains 18 articulated objects drawn from six everyday categories:drawer, lamp, eyeglasses, floppy-disk drive, basket, and phone / laptop stand with three instances per class. For every object we supply the complete set of multi-view RGB-D captures for each recorded interaction state together with the shared camera intrinsics and the per-view extrinsics; a fully reverse-modeling digital twin comprising a USD file that bundles part meshes, textures, physics properties, and joint definitions, a matching lightweight URDF, and watertight PLY meshes for geometry evaluation; and a concise meta.json that enumerates all parts, joint types, axes, limits, and labels every state as either multi-part or single-part motion. This combination of real imagery and fully articulated digital twins enables training, quantitative evaluation, and sim-to-real studies for part-aware geometry reconstruction and joint parameter estimation.
Appendix C Experimental Settings